OpenStudy (anonymous):

What is the simplest form of the expression square root 80

OpenStudy (anonymous):

\( \color{Black}{\Rightarrow \sqrt{80} = \sqrt{2 \times 2 \times 2 \times 2 \times 2 \times 5} }\)

OpenStudy (btaylor):

\[\sqrt{80}=\sqrt{2\times2\times2\times2\times5}\] If a factor is inside twice, you can take it out. So here, you can take out two twos. so you get \[2 \times 2 \times \sqrt{5} = 4\sqrt{5}\] @SirIsaacNewton , you have one too many 2 factors.